# Designing Beyond Language: Sociotechnical Barriers in AI Health Technologies for Limited English Proficiency

### v1

@@ 4.3. Reading and digital literacy limited technology use (RQ1)
In some cases, navigators aimed to circumvent reading literacy issues via technology accessibility features such as sending audio messages rather than text or touching pictures to navigate (P6). However, these practices sometimes collided with many clients’ lack of digital literacy. P6 described an unexpected challenge when accommodating Guatemalan families with low reading literacy:
@@
“I created a flyer with pictures and a QR code that they could scan and hear the information in Q’anjob’al, but we learned that a lot of families didn’t know how to use a QR code. I thought this was a really easy way for them to be able to hear information in their native language, but then there was another barrier.”
@@
Digital literacy varied greatly across LEP individuals. Navigators attributed generational differences to the range of familiarity with technology: “Most adults never had to use a computer before. Families here for many years already have teenage or adult children who handle technology, so adults don’t feel the need…even to use email. However, new immigrants, here for 5-10 years, force themselves to learn technology because even job applications are online now” (P10). P8 states that most of his adult clients still use flip phones or landlines and do not feel the need to use a smartphone, even if given one. On the other hand, navigators that worked with teenagers and young adults explained that their younger clients are comfortable with technology. P10, a school district family liaison, explained that it is easy for teenagers that had never used a phone or computer before immigrating to the U.S. to adopt these technologies effectively.
@@
Low digital literacy often stemmed from barriers to access technology in the first place, creating additional sources of stress and confusion when seeking medical care. Clients of P13, a social worker in a major metropolitan area, often ”use old phones…which are now considered obsolete. They run into connection problems, apps that don’t work well, or just really slow devices. Patients sometimes give up. That adds stress - on top of being sick, now they’re also struggling with technology.” P3 echoed this sentiment, citing financial burdens as a barrier to accessing usable phones or apps. Several navigators said very few of their clients owned laptops or computers, and P6 and P14 said most of their clients did not have stable internet at home (P6’s clients lived in trailer parks and had to go outside the trailer to get better phone service) and instead relied on mobile hotspots. P4 noticed that most individuals do not even know about new technology, highlighting a lack of educational resources in the community. This unfamiliarity with technology creates discomfort: P12 described that one of the students in her digital literacy class ”was so nervous, scared to even touch the computer. She didn’t want to break it.” Barriers to access and low digital literacy caused patients to struggle with even small tasks that seemed trivial even to navigators (P6). For example, patient navigators needed to show their clients how to join Zoom meetings (P6), maneuver the cursor on computers (P12), and overcome frustration with applications that require verification or extra steps (P13).
@@ 4.4. Privacy concerns further discourage use of technology (RQ1)
Interestingly, navigators drew distinctions between types of data that their clients would be comfortable sharing with clinics and technology. P14 outlined the difference as: ”They’re very protective of sensitive information like Social Security numbers or bank accounts. But they’re comfortable sharing details about feelings, medical issues, recipes, or business ideas—things that can’t be stolen or hacked.” Navigators generally agreed, drawing the line between ”willing to give symptoms and medical history and maybe an email” but ”hesitant with phone numbers, addresses, or other personally identifying information” (P6). These boundaries also extended to AI, as most navigators reasoned that clients would be comfortable sharing their symptoms with AI to get information, but would be less willing to log in to an account and save their chat history.
Navigators employed various strategies to assuage their clients’ privacy concerns. P4 emphasized to her clients the confidential nature of their interactions, informing them of their rights to take action or bring her to justice if their information is shared with any outside sources. P3, P8, P10, P12, and P13 employed similar strategies but also noted the trade-offs: as P3 explained, ”We need to sometimes reassure them that…the state won’t share their information. The forms even say the information is confidential. Sometimes we tell them they’re not required to give certain information, so many clients end up providing only very basic information.” P10 emphasizes that her clients need reassurance that their information would not be shared with anyone else, leveraging the personal relationship and trust she established with her clients in order to obtain their information to support them.
@@ 4.5. Opportunities for AI: reducing social barriers and institutional constraints (RQ2)
Navigators identified two ways that AI systems would be uniquely positioned to support LEP individuals seeking care: reducing sentiments of alienation and alleviating resource constraints.
Navigators described that many of their clients felt unheard, rejected, or embarrassed by the healthcare system. P1 explains that ”in Hispanic culture, people like to talk, share details, and give background. Appointments are usually rushed, so doctors usually want ’yes/no’ answers about symptoms, but patients may tell their whole life story first. That mismatch in communication style can definitely be a barrier.” Other patients left appointments feeling like their concerns went unheard due to language or cultural misunderstandings that impeded their ability and confidence to advocate for themselves (P5, P11). Oftentimes institutional restrictions on interpreters prevented them from addressing these issues. At P6’s clinic, the interpreters must strictly provide word-for-word translations, barred from intervening when they perceive misunderstandings between patients and providers. Patients also experienced embarrassment or discomfort when describing medical conditions, especially if doing so in English, causing the patients to refrain from elaborating on their conditions (P2). Lastly, telehealth methods, albeit convenient, may exacerbate these communication disconnects. P5 lamented that telehealth ”make[s] it hard because you can’t see the patient to gauge their pain, and Wi-Fi issues make it worse.” Collectively, these barriers caused patients to feel alienated and not seek care: ”Because English isn’t their first language, they’re extra cautious. They feel vulnerable and worry about being scammed. They know they’re a group people might take advantage of, so they’re wary” (P14). Patient navigators proposed that some of these issues could be alleviated by AI. For example, P2 reasoned that talking to a chatbot about a medical condition would eliminate the sense of embarrassment or social judgment that resides in a human-to-human conversation. P6 proposed the idea of an AI interface that transcribes patient-provider conversations in real-time and detects potentially confusing concepts or jargon in the conversation, allowing patients to gain clarification in real-time without causing embarrassment by verbally interrupting or admitting their confusion. P11 reasoned that talking to AI as a post-visit debrief could help patients feel better by externalizing their sentiments and receiving comfort.
@@ 4.6. Risks of AI: loss of human connection and validation strategies (RQ2)
”An AI might know the facts, but it’s not Mexican, didn’t grow up in Mexico, so it doesn’t really understand me the way a person would. For example, Mexican remedy traditions—American doctors don’t get them, but other people [from similar cultures] do. AI knows a lot about an egg cleanse, but it hasn’t experienced it, so it doesn’t know why we do it.”
@@
Table 3. Summary of barriers LEP individuals face with respect to health technologies, how those barriers manifest, and associated design implications for AI.:
Barrier Observed Challenges Implications for AI
Cultural differences Traditional or home remedies may conflict with western medical advice Fear of being judged prevents full disclosure Respect for authority discourages contradicting or asking questions Acknowledge cultural practices and values while offering practical steps for care Capture nonverbal cues that signify discomfort Avoid misalignment or inappropriate advice due to lack of cultural context
Literacy Low reading literacy in English and/or Spanish Low digital literacy due to limited experience with usable technology (phones, computers, Internet) Lack of specialized medical knowledge complicates understanding health information Voice- or picture-based controls Capture nonverbal cues that signify discomfort Educational resources and guidance that accompany AI to teach prompting and information validation skills Build simpler models that do not require state-of-the-art devices
Privacy Reluctance to provide personal information Fear that personal data could end up in unwanted entities Hesitancy to engage with digital systems, preferring in-person means Offer options to access AI services without providing personal information; anonymize when possible Minimize storing data and history Guarantee confidentiality and privacy of data
@@ 5. Discussion
Overall, patient navigators in our study were favorable towards the adoption of AI technologies to improve healthcare access and quality for LEP individuals. However, introducing AI into this ecosystem raises questions around how it might reshape communication between stakeholders (e.g., patient-navigator or patient-provider relationships). In this section, we discuss opportunities for developing AI tools for facilitating communication between LEP individuals, providers, and navigators. At the same time, our interviews surfaced a number of risks of AI, and so we consider when it might not make sense to develop new tools. For many patients, trust in a system is inseparable from the policies and institutions that govern it, meaning that even well-designed technologies can fail if they do not address broader systemic inequities. Thus, we discuss the implications of our study within these broader structural contexts. We do not generalize the following discussion and considerations to all LEP individuals or even all Spanish-speaking LEP individuals. Table 3 summarizes our findings of the barriers that LEP individuals face and the associated design considerations for AI tools.
5.1. New AI technology should be developed within current practices and relationships
Navigators discussed the difficulties for LEP individuals to adopt new technologies, skills, or even visit a new clinic. Prior work advocates for integrating telehealth within platforms familiar to the patients, rather than expecting them to adapt to new systems (Anaya et al., 2021). This method consciously and proactively includes marginalized populations. We propose 4 examples of how AI can be integrated according to this principle.
@@ 5.1. New AI technology should be developed within current practices and relationships
First, AI should be deployed on devices and platforms that LEP individuals already use. Navigators explained that their clients were generally more comfortable and had more access to phones (as opposed to computers, laptops, wearables, or other forms of technology). Therefore, AI tools should be mobile-friendly or even mobile-centric. Going further, chatbots could send reminders, answer questions, or translate information along familiar channels, such as SMS or WhatsApp, similar to past health interventions (Unger et al., 2018; Abdulhamid et al., 2024). In these ways, individuals would not be expected to learn to use a new device, download new apps, or navigate complex, unfamiliar interfaces.
Second, for individuals without access to stable internet or the newest technologies, system developers should consider employing smaller (e.g., quantized (Dettmers et al., 2023)) models that run efficiently on low-resource devices or require minimal bandwidth. While such models might sacrifice some compute-intensive capabilities (e.g. deep reasoning (OpenAI, 2025)), allowing patients to access tools without an internet connection could alleviate both digital access issues and privacy concerns. P6’s anecdote of families needing to setup hotspots for mobile trailers to access the internet suggests that having tools that could go where individuals go might be more effective than tools that can do more but only with certain digital access. This ensures that AI tools remain accessible to individuals who may face access barriers with outdated phones or intermittent internet service.
Third, AI systems for LEP individuals should not assume baseline digital or reading literacy (in any language). Navigators described that LEP individuals come from a wide range of education and literacy levels, and how even tasks they thought to be simple and accessible, such as scanning QR codes on a phone, can be obstacles. These anecdotes align with prior work that showed assumptions of baseline literacy or technological familiarity often exclude marginalized users from health technologies or digital services in general (Yao et al., 2022; Alkureishi et al., 2021; Whitehead et al., 2023). Prior work has established that speech-based interactions can overcome literacy barriers (Pradhan et al., 2020). Given the dual threads of agents that can navigate complex digital environments (Shaikh et al., 2025; Mozannar et al., 2025; Huq et al., 2025) and generative voice-based interfaces (Le et al., 2023), there is an opportunity to develop voice-based agents that can bridge both reading and digital literacy barriers. For example, agents might interact with patients as a single voice-based system but help them navigate other tools (e.g., Zoom, digital calendars, and email). Voice-based agents for translation may also serve well in the frequent instances in which interpreters, providers, or navigators converse with patients over the phone.
Lastly, system designers should consider cultural values that may lead individuals to engage with or reject new AI tools. Prior work established that cultural values may influence acceptance and choice of technology (Reinecke and Bernstein, 2011). For instance, navigators emphasized the significance of the respect for authority in Hispanic cultures. In order to encourage comfortable and transparent interactions, chat or voice-based systems should adopt the tone of a navigator or peer, while giving information with the expertise of a provider.
Overall, AI systems should adapt to the circumstances of the users, leveraging their existing skillsets and practices whenever possible and minimizing the amount of new knowledge, technology, or skills they would have to learn. One way to do this is to directly involve patients in the development process (Adus et al., 2023) to lower barriers to incorporating AI seamlessly into their existing practices and cultures.
5.2. Psychology of trust suggests private, smaller models
Adoption of AI in healthcare must account for the psychological realities of LEP individuals, who may hesitate to disclose sensitive health information if they fear it could be misused or linked to immigration status. Experiences of discrimination and marginalization can heighten mistrust in institutions, leading individuals to approach new technologies with caution. This underscores the importance of embedding psychological safety and relational trust into AI systems, rather than focusing narrowly on technical accuracy alone.
However, privacy poses a paradox: while patients are protective of their personal information, tailoring generated responses to their needs requires access to such information. This tension reveals the role of needs assessments, as not all patients want to achieve the same goals or functionalities with AI. Some may only seek real-time translation and may not care as much about their data. Others may want to seek personalized support but feel unsafe giving their personal information to an AI system. P6 suggested features that would allow users to utilize nicknames or other anonymizing information to help patients feel more comfortable without disclosing their data. Respecting patient agency by offering configurable options allows for greater willingness to utilize these technologies. While many chat services have options to not share chat logs, navigators pointed out that their clients would distrust any privacy policy, regardless of any mentioned privacy protections, suggesting that such options might not mitigate concerns. Similar to concerns about digital access, privacy issues suggest that smaller models that can run and store chats locally would be preferable.
@@ 5.2. Psychology of trust suggests private, smaller models
5.3. Designing for the long tail of linguistic and cultural variation
Languages and cultures are not monolithic. In our study, patient navigators highlighted how dialect differences can cause misunderstandings, even if interpretation is available. Furthermore, individuals from predominantly Spanish-speaking countries did not necessarily speak Spanish. Though recent years have seen improvements in the multi-lingual capabilities of the LLMs powering AI (Downey et al., 2024; Huang et al., 2025), our findings suggest linguistic variation can go beyond different languages, dialects, or even across families.
AI systems should prioritize dialect-level adaptation, ensuring that they can flexibly account for regional and cultural nuances. Personalization should extend beyond language and into cultural norms, values, and practices (Liu et al., 2025a). For example, AI systems should acknowledge how cultural beliefs around medicine intersect with biomedical advice. AI systems should also capture nonverbal cues in order to fully understand an LEP patient’s perspectives, as patients may not always verbalize their thoughts fully or correctly when under pressure or vulnerability in a setting like receiving medical care. While data might be sparse to train models to do this effectively for every speaker, vision-language models (Deitke et al., 2024) and methods to align models (Shaikh et al., 2024; Wu et al., 2024) to individual preferences might offer an avenue for chatbots to continually learn vocabulary unique to a patient.
Previous studies have shown that providers use online questionnaires, cameras, and wearable technology to collect information about their patients’ personal values and background, in addition to medical information directly related to their symptoms (Berry et al., 2019). AI tools have the potential to not only help dynamically collect information about lived experiences and cultural values, but also provide summaries or explanations to providers of different cultures than their patients by, for example, synthesizing multiple channels of data into insights or points of interest.
5.4. When AI might help, but not be the best answer

### v2

@@ 4.3. Reading and digital literacy limited technology use (RQ1)
In some cases, navigators aimed to circumvent reading literacy issues via technology accessibility features such as sending audio messages rather than text or touching pictures to navigate (P6). However, these practices sometimes collided with clients’ lack of digital literacy. P6 described an unexpected challenge when accommodating Guatemalan families with low reading literacy:
@@
”I created a flyer with pictures and a QR code that they could scan and hear the information in Q’anjob’al, but we learned that a lot of families didn’t know how to use a QR code. I thought this was a really easy way for them to be able to hear information in their native language, but then there was another barrier.”
@@
Digital literacy varied greatly across LEP individuals. Navigators attributed generational differences to the range of familiarity with technology: ”Most adults never had to use a computer before. Families here for many years already have teenage or adult children who handle technology, so adults don’t feel the need…even to use email. However, new immigrants, here for 5-10 years, force themselves to learn technology because even job applications are online now” (P10). P8 stated that most of his adult clients use flip phones or landlines and do not feel the need to use a smartphone, even if given one. On the other hand, navigators that worked with teenagers and young adults explained that their younger clients are comfortable with technology. P10, a school district family liaison, explained that it is easy for teenagers that had never used a phone or computer before immigrating to the U.S. to adopt these technologies effectively.
@@
Low digital literacy often stemmed from barriers to accessing technology in the first place, creating additional sources of stress and confusion when seeking medical care. Clients of P13, a social worker in a major metropolitan area, often ”use old phones…which are now considered obsolete. They run into connection problems, apps that don’t work well, or just really slow devices. Patients sometimes give up. That adds stress - on top of being sick, now they’re also struggling with technology.” P3 echoed this sentiment, citing financial burdens as a barrier to accessing usable phones or apps. Several navigators said very few of their clients owned laptops or computers, and P6 and P14 said most of their clients did not have stable internet at home (P6’s clients lived in trailer parks and had to go outside the trailer to get better phone service) and instead relied on mobile hotspots. P4 noticed that most individuals do not even know about new technology, highlighting a lack of educational resources in the community. This unfamiliarity with technology creates discomfort: P12 described that one of the students in her digital literacy class ”was so nervous, scared to even touch the computer. She didn’t want to break it.” Barriers to access and low digital literacy caused patients to struggle with even small tasks that seemed trivial even to navigators (P6). For example, patient navigators needed to show their clients how to join Zoom meetings (P6), maneuver the cursor on computers (P12), and overcome frustration with applications that require verification or extra steps (P13).
@@ 4.4. Privacy concerns further discourage use of technology (RQ1)
Interestingly, navigators drew distinctions between types of data that their clients would be comfortable sharing with clinics and technology. P14 outlined the difference as: ”They’re very protective of sensitive information like Social Security numbers or bank accounts. But they’re comfortable sharing details about feelings, medical issues, recipes, or business ideas—things that can’t be stolen or hacked.” Other navigators generally agreed, drawing the line between ”willing to give symptoms and medical history and maybe an email” but ”hesitant with phone numbers, addresses, or other personally identifying information” (P6). These boundaries also extended to AI, as most navigators reasoned that clients would be comfortable sharing their symptoms with AI to get information, but would be less willing to log in to an account and save their chat history.
Navigators employed various strategies to assuage their clients’ privacy concerns. P4 emphasized to her clients the confidential nature of their interactions, informing them of their rights to take action or bring her to justice if their information is shared with any outside sources. P3, P8, P10, P12, and P13 employed similar strategies but also noted the trade-offs: as P3 explained, ”We need to sometimes reassure them that…the state won’t share their information. The forms even say the information is confidential. Sometimes we tell them they’re not required to give certain information, so many clients end up providing only very basic information.” P10 emphasizes that her clients need reassurance that their information would not be shared with anyone else, leveraging the personal relationship and trust she established with her clients in order to obtain the information needed to support them.
Table 4. Summary of risks and opportunities for AI health technologies to facilitate communication (RQ2) and associated design implications for AI. Bullet points map observed behavior to a potential implication for AI design.
Theme Observed Behavior Implications for AI
Opportunities Reduce social barriers Patients feel unheard by providers due to language or cultural misunderstandings Patients feel embarrassed or uncomfortable when describing medical conditions Telehealth distance emphasizes alienation Provide on-demand conversational support that eliminates social pressures of a human-human conversation Detect confusion or jargon in real-time conversations and provide simplified explanations Provide on-demand support by allowing patients to externalize their sentiments
mygray      Alleviate resource constraints Long wait times for interpreters Navigators are overburdened to properly follow-up with patients Long work hours increase risk of navigator errors and decreased service quality Provide on-demand translation Proactively send check-in questionnaires to monitor patient progress Automate navigator work, identify errors, and ensure consistency
black Risks      Loss of Human Connection Human presence is significant source of comfort Patients want to feel that someone cares Keep navigators in the loop to use AI alongside patients Use AI to support rather than replace patient-navigator relationships
mygray      Misinformation Patients may not be able to validate AI information or translations Patients may overrely on AI-generated information Build in accuracy checks, safety filters, and plain-language summaries Provide educational supports to teach skills around assessing trustworthiness
black
A table showing 4 rows of opportunities and risks of AI health technologies. Columns include theme, observed behavior, and implications for AI.
@@ 4.5. Opportunities for AI: reducing social barriers and institutional constraints (RQ2)
Having characterized barriers shaping LEP patients’ technology experiences (RQ1), we now describe how navigators envision AI playing a role in communication dynamics and care workflows. Navigators identified two ways that AI systems would be uniquely positioned to support LEP individuals seeking care: reducing sentiments of alienation and alleviating resource constraints.
Navigators described that many of their clients felt unheard, rejected, or embarrassed by the healthcare system. P1 explains that ”in Hispanic culture, people like to talk, share details, and give background. Appointments are usually rushed, so doctors usually want ‘yes/no’ answers about symptoms, but patients may tell their whole life story first. That mismatch in communication style can definitely be a barrier.” Other patients left appointments feeling like their concerns went unheard due to language or cultural misunderstandings that impeded their ability and confidence to advocate for themselves (P5, P11). Oftentimes institutional restrictions on interpreters prevented them from addressing these issues. At P6’s clinic, the interpreters must strictly provide word-for-word translations, barred from intervening when they perceive misunderstandings between patients and providers. Patients also experienced embarrassment or discomfort when describing medical conditions, especially if doing so in English, causing the patients to refrain from elaborating on their conditions (P2). Lastly, telehealth methods, albeit convenient, may exacerbate these communication disconnects. P5 lamented that telehealth ”make[s] it hard because you can’t see the patient to gauge their pain, and Wi-Fi issues make it worse.” Collectively, these barriers caused patients to feel alienated and not seek care: ”Because English isn’t their first language, they’re extra cautious. They feel vulnerable and worry about being scammed. They know they’re a group people might take advantage of, so they’re wary” (P14). Patient navigators proposed that some of these issues could be alleviated by AI. For example, P2 reasoned that talking to a chatbot about a medical condition would eliminate the sense of embarrassment or social judgment that resides in a human-to-human conversation. P6 proposed the idea of an AI interface that transcribes patient-provider conversations in real-time and detects potentially confusing concepts or jargon in the conversation, allowing patients to gain clarification in real-time without causing embarrassment by verbally interrupting or admitting their confusion. P11 reasoned that talking to AI as a post-visit debrief could help patients feel better by externalizing their sentiments and receiving comfort.
@@ 4.6. Risks of AI: loss of human connection and validation strategies (RQ2)
”An AI might know the facts, but it’s not Mexican, didn’t grow up in Mexico, so it doesn’t really understand me the way a person would. For example, Mexican remedy traditions—American doctors don’t get them, but other people [from similar cultures] do. AI knows a lot about an egg cleanse, but it hasn’t experienced it, so it doesn’t really know why we do it.”
@@ 5. Discussion
In this study, patient navigators saw clear potential for AI technologies to improve healthcare access and quality for LEP individuals. However, we avoid framing AI as a solution to structural inequities, as successful introduction of AI depends on how well these systems integrate into these existing interpersonal (e.g., patient-navigator or patient-provider) and sociotechnical ecosystems. Thus, in this section, we discuss opportunities and risks for developing AI tools that facilitate communication between LEP individuals, providers, and navigators. For many patients, trust in and consequent adoption of a system is inseparable from the policies and institutions that govern it (Bahari et al., 2024), meaning that even well-designed technologies can fail if they do not address broader systemic inequities. Thus, we discuss the implications of our study within these broader structural contexts. We do not generalize the following discussion and considerations to all LEP or even all Spanish-speaking LEP individuals.
5.1. Design considerations for embedding AI technology into existing sociotechnical infrastructures
Prior research in care-centered computing advocates for integrating health technologies within existing platforms or systems familiar to patient users  (Anaya et al., 2021; Regenbrecht et al., 2014; Curtis et al., 2025) to minimize burdens and disruptions of adapting to new systems. Consistent with this principle, navigators in our study discussed the difficulties for LEP individuals to adopt new technologies, skills, or even visit a new clinic, as many LEP individuals rely heavily on navigators, family members, or other sources of support to mediate communication, explain processes, or advocate on their behalf. Therefore, rather than demanding users migrate to completely new platforms, we propose integrating AI according to the following principles.
Platform congruence.
@@ 5.1. Design considerations for embedding AI technology into existing sociotechnical infrastructures
To minimize burdens of learning new practices, AI should be deployed on devices and platforms that LEP individuals already use. Navigators explained that their clients were generally more comfortable and had more access to phones (as opposed to computers, laptops, wearables, or other forms of technology). Therefore, AI tools should be mobile-friendly or even mobile-centric. Going further, chatbots could send reminders, answer questions, or translate information along familiar channels, such as SMS or WhatsApp (Unger et al., 2018; Abdulhamid et al., 2024). This ”meet them where they are” approach removes the cognitive load of navigating new interfaces or devices, allowing new AI to integrate quietly in established technological practices.
Small Language Models.
AI must recognize limitations imposed by the digital divide and institutional distrust. Our findings highlight a privacy paradox in which AI requires personal data or third party APIs to tailor generated responses, yet LEP populations are often highly reticent to share their information due to distrust in institutional policies and fearing surveillance. Here, the technical constraints of the digital divide offer a solution to the psychological constraints of trust. By employing either small language models (SLMs) that have fewer parameters or models that can be publicly downloaded (e.g., the Olmo3 model family (Olmo et al., 2025) or gpt-oss (Agarwal et al., 2025)) and run locally, developers can simultaneously address two problems. Models that can run on a mobile device (Li et al., 2024b) can function without stable internet, addressing P6’s observation of families relying on intermittent hotspots in mobile trailers. Psychologically, on-device processing ensures data sovereignty such that because data stays on device, the risk of interception or institutional misuse is minimized (Dubiel et al., 2024). While such models might sacrifice some compute-intensive capabilities (e.g. deep reasoning (OpenAI, 2025)), allowing patients to access tools locally may be more effective than policy statements in reassuring privacy. Furthermore, giving users agency over their data through options to use nicknames, ephemeral chat modes, or privacy-enhancing systems (Pham et al., 2025) can further align the system with user needs for privacy. However, on-device processing does not eliminate all risk (e.g., device loss, coercive access, or organizational policies around tool use), so privacy protections must also include usability-centered safeguards (e.g., quick-delete, no default history, clear threat models). AI design should also explicitly consider risks such as inadvertent disclosure to family members, clinic staff, or authorities; device sharing; and fear of surveillance tied to immigration enforcement. These threats shift design priorities toward minimal data retention, non-identifying access modes, and clear user control over what is stored or shared (Li et al., 2024a).
Enforcing trust and appropriate power dynamics.
Prior work established that cultural values may influence acceptance and choice of technology (Harrington et al., 2022; Reinecke and Bernstein, 2011) due to factors such as alignment of communication style, sensitivity in presenting information, and readability of information (Deva et al., 2025). For instance, navigators emphasized that high respect for authority in Hispanic cultures rendered patients unwilling to fully disclose their symptoms. Therefore, in order to encourage comfortable and transparent interactions, AI systems should adopt the tone of an intermediary between medical authority and accessible peer by integrating computational approaches to politeness (Danescu-Niculescu-Mizil et al., 2013) or style transfer (Reif et al., 2022). In this way, AI can serve as a supportive scaffold that reinforces rather than usurps the established relationships of trust between patients and navigators. A practical pattern is “coach and translator”: the system helps patients rehearse questions in their preferred dialect, flags jargon, and generates respectful question prompts (e.g., “Can you explain that in another way?”) while explicitly encouraging patients to defer medical decisions to clinicians.
@@
5.2. AI-driven personalization to support long tail linguistic, cultural, and literacy variation
Linguistic variation.
Languages and cultures are not monolithic. In our study, navigators highlighted how dialectal differences can cause misunderstandings, even if interpretation is available. Furthermore, individuals from predominantly Spanish-speaking countries did not necessarily speak Spanish. Though recent years have seen improvements in the multi-lingual capabilities of LLMs (Downey et al., 2024; Huang et al., 2025), our findings suggest AI systems should prioritize dialect-level adaptation, ensuring that they can flexibly account for regional, cultural, and individual nuances (Capdevila et al., 2025). However, existing frontier models remain limited in recognizing rich diatopic variability, including in languages like Spanish (Mayor-Rocher et al., 2025; Tellez et al., 2023), highlighting the need for benchmark datasets that can rigorously evaluate LLM capabilities in understanding and responding to diverse dialects (Martínez et al., 2025; Grandury et al., 2025). Personalization here requires systems to align with individual communication preferences, which can be achieved via few-shot learning (Kim and Yang, 2025) or other alignment methods (Shaikh et al., 2024; Wu et al., 2024) that offer avenues for AI to continually learn the patient’s vocabulary and preferences.
Cultural variation.
While health practices are deeply rooted in cultural care routines, navigators explained that patients often conceal cultural practices such as egg cleanses from providers due to fear of judgment or failure to understand. While providers currently use questionnaires and wearable technology (Bharmal et al., 2023; Mak et al., 2025) to capture patient lifestyle and background, raw data alone rarely translates into cultural empathy. AI tools have potential to synthesize culturally-appropriate insights grounded in lived experiences for providers (Peters et al., 2025). One way of achieving this is to pair a generalist high-level LLM and a group of culturally-tuned LLMs in order to incorporate diverse cultural perspectives (Sorensen et al., 2024) and thus validate medical advice against specialized cultural norms. Furthermore, in-context adaption via sociodemographic prompting or cultural role-play (Shaikh et al., 2023; Hwang et al., 2023) can allow the agent to proactively identify and bridge value gaps between patients and providers. In this way, an AI system could, for example, detect hesitance regarding a treatment plan and generate culturally relevant analogies or explanations that align with the patient’s lived experience (Peters et al., 2025). However, this requires a shift from static profiling to dynamic cultural alignment, ensuring that systems avoid stereotyping by continuously learning from the user’s specific feedback and value signals (Fan et al., 2025; Wu et al., 2025a).
Conversely, for patient-facing interactions, AI systems should recognize how cultural beliefs around medicine intersect with biomedical recommendations, framing guidance in a way that is both culturally sensitive and clinically appropriate. Value-sensitive design suggests that systems should acknowledge the user’s lived experiences before pivoting to clinical corrections to maintain trust and adherence (Liao and Wortman Vaughan, 2024). In situations where following cultural beliefs could lead to serious health consequences, such as the case of P10’s student who believed brujería (witchcraft) caused his excruciating leg and spine pain, AI should be governed by context-aware safety alignment (Shen et al., 2025) to gently prioritize biomedical advice, ensuring that critical interventions are not overlooked. AI can minimize the risk of alienating patients during critical moments by framing medical interventions as necessary complements to, rather than replacements for, their cultural practices.
Literacy variation.
Assumptions of baseline literacy or technological familiarity often exclude marginalized users from health technologies or general digital services (Yao et al., 2022; Alkureishi et al., 2021; Whitehead et al., 2023). Literacy, however, is multi-dimensional, encompassing reading, digital, health, and even specialized forms of AI literacy such as scientific and data literacy (Long and Magerko, 2020). Navigators emphasized that LEP individuals often struggle concurrently with multiple dimensions of literacy such that an AI tool that simplifies text to a 5th-grade level is still inaccessible since many patients lack the skills to type queries or navigate interfaces, often relying on other modalities such as voice messages to communicate. Therefore, AI systems should adapt content and modality simultaneously to address multi-dimensional literacy gaps. The complexity of content can be dynamically adjusted by providing definitions of unfamiliar terms, plain-language summaries (August et al., 2023), or simplifications based on a user’s demonstrated reading (Naeem et al., 2025) or health literacy level (Schillinger et al., 2021). Furthermore, AI interfaces for LEP populations should be multimodal by default. Given the dual threads of agents that can navigate complex digital environments (Shaikh et al., 2025; Mozannar et al., 2025; Huq et al., 2025) and generative voice-based interfaces that can overcome reading and typing barriers (Pradhan et al., 2020; Le et al., 2023), there is an opportunity to develop voice-based agents that can, for example, help patients navigate other tools (e.g., Zoom, digital calendars, and email). Voice-based translation agents (Hudelson and Chappuis, 2024) may also support the frequent over-the-phone interactions between interpreters, patients, and providers. Furthermore, text-to-image models can generate visual explanations (Wu et al., 2025b) such as diagrams showing how to take medication, tailored to the specific medication or dosage. Leveraging generative AI capabilities to move beyond text-centric interventions can help bridge the multifaceted literacy gaps faced by LEP individuals.
5.3. When AI might not be the best approach